### v3

# Adversarial Evasion Attacks Practicality in Networks: Testing the Impact of Dynamic Learning Authors’ draft for soliciting feedback: April 4, 2024

Abstract
Machine Learning (ML) has become ubiquitous, and its deployment in Network Intrusion Detection Systems (NIDS) is inevitable due to its automated nature and high accuracy compared to traditional models in processing and classifying large volumes of data. However, ML has been found to have several flaws, most importantly, adversarial attacks, which aim to trick ML models into producing faulty predictions. While most adversarial attack research focuses on computer vision datasets, recent studies have explored the suitability of these attacks against ML-based network security entities, especially NIDS, due to the wide difference between different domains regarding the generation of adversarial attacks.
To further explore the practicality of adversarial attacks against ML-based NIDS in-depth, this paper presents three distinct contributions: identifying numerous practicality issues for evasion adversarial attacks on ML-NIDS using an attack tree threat model, introducing a taxonomy of practicality issues associated with adversarial attacks against ML-based NIDS, and investigating how the dynamicity of some real-world ML models affects adversarial attacks against NIDS. Our experiments indicate that continuous re-training, even without adversarial training, can reduce the effectiveness of adversarial attacks. While adversarial attacks can compromise ML-based NIDSs, our aim is to highlight the significant gap between research and real-world practicality in this domain, warranting attention.
Index Terms:
machine learning, dynamic machine learning, adversarial attacks, network intrusion detection systems
I Introduction
Machine Learning (ML) applications are present in every corner. ML found its way to network security applications due to its (1) automated nature and (2) ability to process and classify high volumes of data with high accuracy. Thus, ML deployment in Network Intrusion Detection Systems (NIDS) is a foregone conclusion. However, in recent years, researchers highlighted that ML has a lot of drawbacks and flaws. To name a few, ML has privacy concerns [1] and suffers from reproducibility crises (when the results of a model can not be replicated using the exact code and dataset) [2]. Nevertheless, one ML challenge shined the most in the research community; adversarial attacks. In a nutshell, these adversarial attacks aim to trick ML models, making them produce faulty predictions. There is no doubt that these attacks are a menace to ML. However, it was observed that the majority of adversarial attack research uses computer vision datasets to evaluate its severity [3]. Thus, recently, a few researchers [4] [5] [6] have explored the practicality of adversarial attacks against ML-based network security entities, especially Network Intrusion Detection Systems (NIDS), and implied that attacking them is much more complicated than computer vision.
Contributions: In this paper, we study the practicality of evasion adversarial attacks against NIDS and make three distinct contributions; First, we identify the questionable practicality prerequisites required to carry out evasion adversarial attacks against NIDS. We introduce an attack tree, a threat modeling technique, to visualize various scenarios attackers might use to launch evasion adversarial attacks against ML-NIDS. The attack tree helps identify leaf nodes with questionable feasibility. Secondly, we extend the research discussion on the practicality of adversarial attacks against NIDS by introducing a taxonomy of practicality issues associated with adversarial attacks against ML-based NIDS. This taxonomy includes issues reported in the literature, along with our own questions and observations. The third contribution emerged from our observation that the overwhelming bulk of adversarial attack research deals with ML approaches as static models. However, nowadays, numerous ML models are dynamic and constantly evolving. Approaches like stream learning [7] or online machine learning [8] constantly update their models to make them adaptive to new cases and data (prevent concept drifts). Therefore, we explore and test the effect of continuous training on adversarial attacks against NIDS. To the best of our knowledge, these are the first experiments on the impact of continuous training on such invasions against ML-based NIDS. This paper builds upon the preliminary findings presented in our poster [6].
The rest of the paper is organized as follows: Section II delves into the relevant background. In Section III, we review previous research and surveys. The attack tree, which shows the various techniques attackers might employ to launch evasion-based adversarial attacks against ML-NIDS, is presented in Section IV. Section V introduces the taxonomy of practical challenges and issues related to adversarial attacks against ML-NIDS. Our investigation environment and methodology are presented in Section VI. The outcomes of our study, illustrating the impact of continuous training on adversarial attacks in the NIDS domain, are outlined in Section VII. Section VIII encompasses discussions and lessons learned from our research. Finally, Section IX concludes our work and provides insights into future directions.
II Background
Figure 1: The Deployment of Network Intrusion Detection System
II-A ML-based NIDS
Network intrusion detection systems (NIDS) are tools designed to identify threats against a network. They send alerts to network management in case of breaches. Figure 1 shows the deployment of Network Intrusion Detection Systems. There are two main NIDS techniques: signature-based and anomaly-based. Signature-based detection identifies threats by detecting predetermined attack patterns (signatures). This type of detection has a low false positive rate in an environment of well-defined attacks. However, its main disadvantage is that it cannot detect newer or unknown attacks [9]. Anomaly-based detection, on the other hand, identifies attacks by detecting deviations from normal activity. This approach makes it able to detect unknown attacks. However, because of the difficulty of defining the boundary between normal and abnormal behavior, anomaly-based detection can have a high false positive rate [10]. Therefore, to solve the accuracy issues in anomaly-based detection, researchers have explored the use of machine learning approaches (like Support Vector Machine (SVM), Decision Tree (DT), Artificial Neural Networks (ANN), etc.) [11] and found them very accurate in classifying large amounts of data. However, some researchers highlighted that machine learning can be vulnerable to adversarial attacks [1]. We will further explore these attacks later in this section.
II-B Dynamic Machine Learning
Concept drift in machine learning is the unforeseeable statistical changes in the prediction target’s properties [12]. In other words, the norm changes over time, so the ML model makes wrong predictions. In an adversary environment, where attackers change their techniques all the time, like the network/cybersecurity domain, concept drifts will more than likely occur [7]. Thus, to counter these drifts ML models must be dynamic, i.e, continuously evolving. Numerous terms are used to describe dynamic ML models (e.g., Online ML, Evolving ML, Continuous/Continual ML, and Stream ML). In the ML community, the implementation and definition of these kinds of ML techniques could differ from one study to another. However, the following concept is common; "The initial training set loses its validity as time passes due to changes of conditions in the aimed task. Thus, a mechanism to update a given model in order to adapt to new conditions is necessary" [8]. In simple terms, dynamic ML models must be updated and re-trained all the time to resist the effect of concept drift. There are different techniques for re-training (or continuous training); for example, adaptive windowing, where the model is re-trained when a concept drift is detected [13]. However, in practice, there are more elementary solutions, with less overhead on the system, like periodically re-training the model with recent data. In our tests, we use the latter approach.
II-C Adversarial Attacks Against Network Security and Defenses
Adversarial examples, first discovered in 2013 by Szegedy et al. [14], are inputs that cause machine learning models to make incorrect predictions despite being very similar to legitimate examples. Soon after, Goodfellow et al. [15] introduced the FGSM (Fast Gradient Sign Method) attack, which creates adversarial examples by adding small perturbations equal to the sign of the elements of the gradient of the cost function. These attacks have become more sophisticated and diverse. In the next subsections, we will explore their different types and categories.
II-C1 Taxonomy of Adversarial Attacks in Network Security
Attackers’ Knowledge:
In White-box Attacks, the attackers have complete knowledge of the attacked ML model (which ML algorithm was used, gradient, hyperparameters, etc.). The complete opposite is Black-box Attacks, where the attackers have no knowledge of the attacked ML model. However, in most cases, researchers assume that the attackers crafting black-box attacks have access to the target ML system’s output. Thus, the attackers can create a substitute model (pretty close to the attacked model) by querying the targeted system. Then they use that substitute system to get the knowledge they need to attack. Grey-box Attacks lie between them (the attackers have "some" knowledge of the attacked ML system).
Attackers’ Approaches:
In Evasion Attacks, which is the main focus of this paper, attackers aim to deceive the ML model during the decision-making process by adding carefully crafted perturbations to the input data. Like [16], where a perturbation is added to the features to make the IoT ML intrusion detection model output the wrong classification. As for Poisoning Attacks, the attackers add malicious samples to the training datasets (dataset contamination). Like [17], where a data poisoning attack against a federated learning-based IoT intrusion detection system allowed the adversaries to implant a backdoor into the ML detection model to wrongly classify malicious traffic as benign. In Model Stealing Attacks, an attacker can probe a black box model to extract all the vital information to create a "clone" model very close to the target model. It is worth mentioning that ML models in a monumental system like NIDS are very expensive, and requires a lot of resources to create. Thus, stealing them will be a massive loss to the associations that created them. The main focus of this paper is evasion attacks. Regarding Backdoor Attacks, the attackers aim to embed a backdoor into the ML model. The backdoored model operates normally when processing inputs lacking the trigger. However, it is misdirected to perform the attacker’s sub-task once the secret trigger is presented in the input [18].
Attack’s Space:
Okunle Ibitoye et al. [19] presented the space term in adversarial attacks taxonomy in the network security domain. In Feature-space attacks, the attacker only modifies (by adding perturbations) the ML model input feature vector to craft the attack. As for the Problem-space attacks, the attacker modifies the actual file (malware, pcap, PE (Portable Executable), etc.) to trick the attacked model.
II-C2 Defenses Against Adversarial Attacks
There are multiple approaches to defending ML models against adversarial attacks. Here are a couple of examples: Adversarial Training: This involves inserting adversarial examples into the training dataset to enhance the model’s robustness against such attacks [20] [21]. Feature Reduction: Reducing the feature sets may decrease the effectiveness of perturbations, thereby making the ML classifier less vulnerable to evasion attacks. This technique has shown promising results in ML spam and malware detection models [22].
(a) Adversarial Example Generation in the Computer Vision Domain
(b) Adversarial Example Generation in the Network Security Domain
Figure 2: Adversarial Examples Generation
III Related Work
In this section, we review previous related work and classify them into two groups; adversarial attacks against ML-NIDS, and the practicality of adversarial attacks against ML-NIDS.
III-A Adversarial attacks against ML-NIDS
Numerous researchers have examined adversarial attacks against machine learning-based network intrusion detection systems. Alatwi et al. [23] surveyed and categorized the studies of adversarial attacks against ML-based NIDSs and briefly discussed their applicability in real-world scenarios. He et al. [24] has also conducted a comprehensive exploration of white- and black-box adversarial attacks on DNNs, and their potential relevance in NIDS. They also analyzed existing defense mechanisms against these attacks. Furthermore, Martins et al. [25] examined adversarial machine learning against intrusion and malware detection systems. They concluded that a wide variety of the reviewed attacks could be effective, however, their practicality was not been tested enough in intrusion scenarios. Furthermore, They have stated that some of the datasets in intrusion scenarios are outdated, and most of them are similar and lack variety.
III-B Practicality of Adversarial attacks against ML-NIDS
Some studies have focused on the practicality of adversarial attacks in the network intrusion detection domain; Vitorino et al. [26] presented an overview of adversarial learning techniques capable of generating realistic examples to deceive NIDSs. Apruzzese et al. [4] aimed to specify and model the abilities and circumstances necessary for attackers to execute practical adversarial attacks on network intrusion detection systems. Both [26] and [4] primarily reviewed and evaluated attacks from research literature but did not consider the dynamic nature of modern ML models. In a separate study, Apruzzese et al. [3] emphasized the existence of a gap between researchers and practitioners in the field of adversarial attacks. Their primary focus was not on NIDSs; however, through collaboration between industry and academia, they articulated the general viewpoint that "Real attackers don’t compute gradients." In other words, they were unable to identify an attacker who computes gradients to craft real-world adversarial attacks.
Merzouk et al. [5] attempted to inspect the practicality of adversarial attacks to evade network intrusion detection systems. They highlighted some limitations of adversarial attacks against ML-NIDS. However, they have only concentrated on the feature space attacks.
To the best of our knowledge, none of the previous work has explored and tested the practicality of adversarial attacks against dynamic ML models in the intrusion detection domain.
{forest}
for tree= draw, text centered [ Adversarial
Evasion Attack
Against ML-NIDS [Create Adversarial
Packets that produces
perturbed features
(Problem-Space), angle below [Access Network
Packets, name =y] [Add perturbations
on Network Packets, angle below [Maintain Network
Functionality] [Maintain Malicious
Functionality
?] [Reverse Engineer
the packets in
a way that
produces perturbed
features after
the features
extraction, angle below [Pick perturbable
Features[Predict some
Feature Vector Items] [Insider
Knowledge
(White-Box)
?]][Know
up-to-date
information of
the model, name =x [Insider
Knowledge
(White-Box)
?] [Query
the model
(Black-Box) [Access Model’s
input and
output
?]]]]]] [Adversarial
Features
(Feature-Space), angle below [Access Feature
Vector
?] [Add perturbations
on Perturable
Features, angle below[Pick perturbable
Features[Predict some
Feature Vector
Items] [Insider
Knowledge
(White-Box)
?]] [Know
up-to-date
information of
the model [Insider
Knowledge
(White-box)
?] [Query
the model
(Black-Box) [Access Model’s
input and
output
[…]
IV Identifying Infeasible Adversarial Attack Prerequisites Through an Attack Tree Threat Model
The majority of the adversarial attacks were designed and tested in the literature using computer vision datasets [3]. Because of the focus on the computer vision datasets, it is not clear if these attacks will be applicable in the same way to the network/cybersecurity domain which has more adversaries and more security constraints and measures, making it much more complicated to evade. This contrast is depicted in Figure 2, which illustrates the crafting of evasion adversarial attacks in both the computer vision and network domains. Thus, in this section, we present an attack tree, depicted in Figure 3, that showcases the various strategies attackers might employ to execute evasion adversarial attacks. Additionally, we identify leaf nodes (prerequisites) with questionable feasibility marked by the ’?’ sign.
The attack tree presented in Figure 3 outlines an Adversarial Evasion Attack strategy targeting Machine Learning-based Network Intrusion Detection Systems (ML-NIDS).
The tree diverges into two main branches, representing the two alternative ways attackers can fool the model: Adversarial Features (Feature space) and Adversarial Packets (Problem-Space). The right branch focuses on the manipulation of feature vectors to fool ML-NIDS. To achieve this, attackers require access to the feature vector and the ability to introduce perturbations to specific features. However, directly accessing the feature vector presents a significant obstacle for external attackers, rendering this approach highly improbable. On the other hand, introducing perturbations to the feature vector necessitates identifying perturbable features (not all features are susceptible to manipulation, as will be discussed further) and requires up-to-date knowledge of the underlying ML-NIDS model. Identifying perturbable features necessitates information about the specific features employed by the target ML-NIDS. Acquiring this information through insider knowledge is highly unlikely. An alternative option for the attackers is to make informed assumptions or predictions regarding certain features, an approach subject to uncertainties. As for up-to-date knowledge of the model, the attacker could get it by insider knowledge (questionable practicality) or by querying the model. However, as illustrated in Figure 1, accessing the model’s input and output is nearly impossible for an outsider. This presents a significant challenge for attackers attempting to launch adversarial attacks against ML-NIDS.
In the left branch of the tree, we delve into Problem-Space attacks, which involve manipulating network packets to evade ML-NIDS detection. In this scenario, the attacker needs to gain access to network packets and introduce perturbations to them. Accessing network packets is comparatively more feasible than accessing the feature vector of the model, making problem-space attacks appear more practical than feature-space attacks. However, adding perturbations to network packets requires the attacker to preserve network functionality, maintain malicious functionality (since it’s an attack), and reverse engineer the packets to produce perturbed features after extraction. We find maintaining the malicious functionality of network packets presents an uncertain feasibility, as existing adversarial attacks against ML-NIDS lack a comprehensive evaluation of the maliciousness of their generated examples. These attacks primarily focus on evading detection by the model, neglecting the preservation of the intended malicious behavior [24]. Additionally, the process of reverse engineering the packets entails similar requirements as the feature-space attack outlined in the right branch of the tree.
The attack tree illustrates that achieving the root node goal is highly challenging. Although problem-space attacks may seem more practical, we find that both problem and feature-space attacks have limited suitability in real-world scenarios.
{forest}
for tree= draw, text centered [ Practicality Issues of Adversarial
Evasion Attacks Against ML-NIDS [Attackers’
Knowledge[White-Box [Own
ML-NIDS [Very unlikely
for an
attacker to
get knowledge
of the model]] [Bought
Pre-built
ML-NIDS [Different settings
at deployment
still Very
unlikely
for an attacker
to get knowledge
of the model]]] [Black-Box [Basic
Knowledge][Querying
NIDSs]]] [Attack’s
Space [ Feature
Space
Perturbations [Accessing the
Feature Vector] [NIDS
Features ,name=x [Imperturbable
Features] [Correlated
Features] [Heterogeneous
Features] [Few remaining
Perturbable
Features]]] [ Problem
Space
Perturbations, name =y [Network
Functionality][Malicious
Functionality][Inverse
Feature-
Mapping
Problem] [Side-
Effect
Features]]] [Dynamic
ML [Attacker
must
always have
up-to-date
knowledge
of the model
which is
[…]
Figure 4: Taxonomy of Practicality Isues of Adversarial Attacks Against ML-NIDS, Directed Acyclic Graph (DAG)
V Taxonomy of Practicality Isues of Adversarial Attacks Against ML-NIDS
In this section, we explore the issues and challenges of adversarial attacks against ML-NIDS. Our taxonomy, illustrated in Figure 4, highlights the gap between research and real-world scenarios. It offers a different perspective on the practicality of adversarial evasion attacks against NIDS compared to the attack tree by summarizing most reported practicality issues from the literature and our notes within a single Directed Acyclic Graph (DAG). This taxonomy aims to bridge the gap between research and real-world scenarios.
V-A Attackers’ Knowledge
V-A1 White-Box Feasibility
Organizations with dedicated security teams may possess the expertise and budget to develop their own Network Intrusion Detection Systems (NIDS). Alternatively, other organizations may choose to purchase pre-built solutions. In the former scenario, acquiring knowledge of a private model is considered highly improbable for an attacker, as white-box attacks require extensive information such as the ML model utilized, gradients, feature set, and more. Even with publicly available pre-purchased models, the diverse deployment settings make obtaining accurate knowledge challenging for an attacker.
V-A2 Black-Box Feasibility
As mentioned in section II, attackers in black-box settings have no knowledge of the target ML model. There is no doubt that black-box evasions are the more practical type of adversarial attack [23] as obtaining all the needed information about the target model is extremely difficult in the network security domain. In the next sub-section, we will explore the feasibility of black-box attacks against ML-based NIDS.
Basic Knowledge
In recent years, researchers successfully crafted adversarial examples without previous knowledge of the target’s learning algorithm, gradients, parameters, training data, or hyper-parameters [27]. However, only a minority of these researches use intrusion, malware, or spam detection datasets. As seen in figure 2a, to generate adversarial examples in the computer vision domain, the attackers adds perturbation by slightly changing the pixels’ values. The medium in the computer vision domain is well-known as the attackers are perturbing a digital image that is visible to them. On the other hand, in network intrusion, the attackers don’t know the medium they are trying to perturb. Each ML-based network intrusion detection system uses its unique pre-processing/feature extraction and selection techniques. If the attackers don’t know the selected features and the pre-processing methods of the target model, then they don’t know what to perturb. Kuppa et al. [28] labeled their attack as black-box, however, they assumed that the adversaries have previous knowledge of the target model’s feature set, feature extractor, and feature transformers. We found that most of the "black-box" attacks in the NIDS domain make similar assumptions. In our opinion, if the attackers know all this information about the features management and selection of the target model, their attack cannot be considered black-box.
Querying NIDS
As mentioned in section II, one of the main corners of black-box adversarial attacks is querying the target model to obtain a substitute model. Then using that alternative model, the attackers can get all the knowledge they need to craft their adversarial attack. It would be reasonable to assume that only the system management teams have access to the output of NIDS, as seen in Figure 1. Thus, it is not clear how querying NIDSs in black-box attack settings is possible if the attacker is not an insider.
V-B Attack’s Space
V-B1 Feature-Space Perturbations
In feature-space attacks, the attacker just changes the ML model input feature vector to fool the target model by adding perturbations to the features.
Accessing the Feature Vector:
Normally, attackers don’t have access to the feature vector after the feature extraction phase in the ML-based NIDS pipeline. So how would feature-space adversarial attacks against NIDS be considered practical? That is an important question to answer considering that most of NIDS’s adversarial attack research is in feature-space attacks [4].
NIDS Features:
As previously mentioned, knowing which features, and feature transformation techniques are used by the target model is a very hard task for the attacker. However, even if the attackers know the selected features and have access to them, adding perturbations to the ML-based NIDS features is a complicated task [5]. Unlike the computer vision domain, numerous constraints make crafting adversarial attacks in the NIDS space tough.
  • Imperturbable Features: A sizable portion of features in the NIDS datasets are boolean (such as flags) and discrete numbers (such as the protocol feature). The presence of these features doesn’t give flexibility to the attackers to add perturbations that maximizes the loss to fool the target model. For example, the attacker cannot add a perturbation to a flag in the positive direction if its value is 1 (because the maximum value is 1). The same goes for a discrete feature, for instance, the attacker cannot add an arbitrary value to the protocol feature because some protocol numbers are unassigned (e.g., unassigned IP address). Moreover, some numerical values have similar limitations, for example, the attacker cannot change the packet length feature to a negative value.
  • Correlated Features: NIDS typically extract various features from network traffic to identify malicious activity. These features are not always independent and can exhibit dependencies or correlations between them. For example, TCP flags in packets and TCP as the transport protocol are correlated [29]. Unlike in computer vision, in the NIDS domain, attackers cannot simply add perturbations to independent pixels; instead, they must consider the complex relationships between different features, which complicates the crafting of attacks in the NIDS domain.
  • Heterogeneous Features: In the NIDS domain, features extracted from network traffic fall into different categories with varying types and characteristics. This diversity poses challenges for attackers crafting adversarial attacks. Unlike computer vision, where the perturbed medium is of one type, the features in the NIDS domain are diverse, including numerical, categorical, ordinal, textual, etc. Adding perturbations to these various types of data is not as straightforward as adding perturbations to pixels. Moreover, even features with the same type may have different characteristics; for example, the range of one numerical feature might be completely different from another numerical feature, adding another layer of complexity to crafting the attack.
  • Few Remaining Perturbable Features: As mentioned in Section II, feature reduction can be a potential defense strategy against adversarial attacks, as it reduces the attack surface and increases the model’s robustness. Due to the NIDS feature constraints mentioned in this section, the NIDS domain has fewer perturbable features compared to other domains like computer vision, which act as a natural line of defense against adversarial attacks.
V-C Problem-Space Perturbations
In problem-space adversarial attacks, the attacker modifies the actual packet to deceive the target model. These attacks are much more practical than feature-space attacks, as the attacker only has access to the file and not the feature vector. However, adding perturbations to network packets is extremely difficult. Usually, a problem-space attack begins with feature-space perturbations. Then, these feature-space manipulations are translated into real-world modifications of network packets, a process referred to as the Inverse Feature-Mapping Problem [30]. This problem poses a significant barrier for attackers and often complicates, and in most cases, prevents, the direct application of gradient-based feature-space attacks to discover problem-space adversarial examples [30]. Moreover, there is no guarantee that an optimal solution in the problem space will be similar to the target adversarial features [24].
Furthermore, the inverse feature process mapping can introduce unintended consequences known as side effect features [30]. Side effect features refer to the unintended characteristics or features that emerge when applying adversarial perturbations to data. These features arise due to the correlations, limitations, and constraints involved in translating manipulations from the feature space to real-world modifications in the problem space. In simpler terms, side effect features are altered features that arise just to satisfy the constraints of the problem space. It has been observed that these features do not follow any particular direction of the gradient, meaning they can have both positive and negative effects.
On another note, the attacker must know which features and feature transformation techniques are utilized by the target model so that the modification of the packet produces perturbations to the features after the feature extraction phase. This requirement means that problem-space attacks inherit all the NIDS feature constraints discussed in the previous subsection from feature-space attacks, as the attacker must address all of them.
Furthermore, even if the attackers succeed in modifying the packet correctly to trick the model, these modifications will more than likely alter the network and malicious functionality of the perturbed packets [31]. For example, changing flags or length values in the header will render them inaccurate, and packets with faulty headers are immediately discarded at the destination, resulting in a loss of network functionality. Moreover, adding perturbations to the payload of the packet could change its intended malicious functionality. For instance, if the attackers aim for an SQL injection attack, modifying the packet’s content will most likely nullify its hostile purpose. It is worth mentioning that existing problem-space adversarial attacks against ML-NIDS lack evaluation of whether their generated examples maintain malicious functionality [24]. This means current attacks cannot guarantee that the altered packets will still be harmful.
Therefore, as depicted in Figure 2, crafting adversarial attacks against NIDS is much more complicated compared to computer vision. While adding small perturbations to pixels in computer vision may not change their functionality, in the case of NIDS, such attacks can potentially alter both their malicious behavior and network functionality, as shown in Figure 2b.
Figure 5: Attacking Scenario with Continuous Training: the impact of adversarial attacks before (attacking in Day n) and after re-training (attacking in Day n+1 and Day n+2)
V-D Adaptivity to Dynamic ML
The vast majority of adversarial attacks depend on the gradients of the target models to maximize the loss and make the attacked model produce faulty decisions. On the other hand, as previously mentioned, there is a trend to use dynamic ML which is reasonable to assume that NIDSs would adopt. Because the gradients in dynamic ML are always changing, that could add to the complexity of creating an attack. To test this assumption, we investigated the effect of dynamic learning on adversarial attacks against ML-NIDS. In the next sections, we will discuss our experimental environment and results.
VI Investigating of The Effect of Dynamic Learning on Adversarial Attacks Against ML-NIDS
VI-A Overview
As noted in Section V-D, we believe that defence that adopt dynamic model should reduce the effectiveness of adversarial attack. We set up and conduct experiments to test the effect of continuous training on the effectiveness of adversarial attacks against ML-NIDS. As illustrated in Figure 5, our experiments test the impact of adversarial attacks before (attacking on Day n) and after re-training (attacking on Day n+1, and Day n+2). In this section, we will explore the ML model, dataset, metrics, and continuous training criteria used to test the practicality of adversarial attacks against dynamic ML-based NIDS.
VI-B Dataset
The CSE-CIC-IDS2018 [32] dataset was used in our experiments. Although only 12 percent of the network intrusion detection research utilizes this dataset, it was found that it is currently the most up-to-date, and ML models have better accuracy when it is used in training and testing [11]. Moreover, the CSE-CIC-IDS2018 dataset has a low number of duplicate and uncertain data [33]. The dataset was created by Communications Security Establishment (CSE) and the Canadian Institute for Cybersecurity (CIC) in 2018/2019. In the CSV format files, each tuple has 80 features. During the creation of the dataset, numerous scenarios were constructed, and data were collected and edited daily. The dataset contains six different classes (labels); Benign, Bot, Brute Force, DoS, Infiltration, and SQL Injection. However, the authors of [34] found that CSE-CIC-IDS-2018 suffers from numerous issues regarding feature generation and labeling. We agree with their findings as we discovered a considerable number of outlier samples in the dataset that need cleansing before the standardization and training processes for constructing more robust ML models.
Figure 6: Target ML-based NIDS
VI-C The Target ML-based NIDS
As seen in Figure 6, the target ML-based NIDS consists of multiple phases. We crafted our own ML-NIDS for our experiments. However, we didn’t create a feature extraction stage, as we used the features provided by the CSV files of the CSE-CIC-IDS2018 dataset. We will explore our ML-NIDS in the next subsections.
VI-C1 Pre-Processing
The pre-processing steps employed in our experiments contained several key procedures.
Initially, a process of data cleansing was implemented, whereby features containing nulls, infinite values, NaN (Not a Number), or single values were removed. Additionally, thousands of numbers in the dataset were stored in string formats, which had to be converted back into numeric values.
Moreover, we found numerous outlier samples, which are values that deviate significantly from the rest of the dataset. These atypical data points can present challenges for statistical analyses, like mean and standard deviation calculations, which are essential for standardization. Furthermore, in ML training datasets, outliers can disrupt model performance, mislead the training process, lead to non-generalizable models, and increase variance, potentially causing bias and reduced accuracy [35]. For example, as seen in Figure 7a, a minuscule number of records (4 samples out of millions) of "Bwd Pkt Len Max" Feature are in the neighborhood of 6000, which is extremely far from the rest of the values. Therefore, we created a method to exclude these outlier samples from the dataset. This technique detects and removes samples with a probability of less than 0.0005 in the cumulative distribution function and those with a probability of more than 0.9995, but only if the number of these records is less than a threshold. The Frequency of "Bwd Pkt Len Max" Feature values Before and After Data Cleansing can be seen in Figure 7. After excluding the small number of outliers, the range of values for the feature changed from 0 to 6000 to 0 to 1500, as seen in Figure 7b. This outlier-excluding method is applied to all numerical features. Afterward, We performed standardization and one-hot key encoding on the dataset to prepare it for training machine learning models.
(a) The Frequency of "Bwd Pkt Len Max" Feature Values Before Data Cleansing
(b) The Frequency of "Bwd Pkt Len Max" Feature Values After Data Cleansing
Figure 7: The Frequency of "Bwd Pkt Len Max" Feature Values Before and After Data Cleansing
Subsequently, as seen in Figure 8,we conducted a correlation analysis to identify highly correlated features, which were then partitioned into mutually exclusive sets to exclude dependent features. Additionally, we calculated the correlation between features and labels and determined the importance order of features using Random Forests on separate decision trees. Finally, based on the obtained results, we selected the most important features for further analysis.
In machine learning, imbalanced datasets can lead to poor accuracy, especially for the minority class [36]. As seen in Figure 9a, the used part of the CSE-CIC-IDS2018 dataset was mildly imbalanced, so we balanced it by using the random under-sampling method [37]. The training dataset’s labels distribution after random under-sampling is shown in Figure 9b.
Figure 8: Dataset’s Features Correlation Matrix
(a) Training Dataset’s Labels Distribution before Random Under-sampling
(b) Training Dataset’s Labels Distribution after Random Under-sampling
Figure 9: Before and After Balancing the Training Dataset
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
(a) FGSM
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
(b) PGD
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
(c) BIM
Figure 10: Accuracy (Y-axis) of the NIDSs before and after the attacks, where Day n represents attacking before re-training, Day n+1 represents attacking one Day after re-training, and Day n+2 represents attacking two days after re-training.
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
(a) FGSM
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
(b) PGD
Pre-attackDay nDay n+1Day n+2000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111ANNCNNSVM
(c) BIM
Figure 11: F1-measure (Y-axis) of the NIDSs before and after the attacks, where Day n represents attacking before re-training, Day n+1 represents attacking one day after re-training, and Day n+2 represents attacking two days after re-training.
VI-C2 Training and Testing
Several ML models were used to create different versions of our experimental NIDS for better comparisons to explore the effect of continuous learning on adversarial attacks.
The first ML model is a fully connected Artificial Neural Network (ANN). The network consists of an input layer containing 50 nodes representing the selected features. Then, two hidden layers with 64 nodes each (using Relu activation function with L2 regularization for the first hidden layer and TANH activation function with L2 regularization for the second hidden layer), which represent the significant attributes describing the nature of the packet, are connected to an output layer with a single node (representing malicious or benign). The ANN uses the ADAM optimizer with an initial learning rate of 0.001. Moreover, the model includes a dropout of 0.4, batch size of 1000, and 20 epochs for training.
Our second model is a Convolution Neural Network (CNN) consisting of three groups of hidden layers. The first group contains a 1D convolution layer with 128 filters, each with a size of 6 and a ReLU activation function. After each convolution layer, we added a dropout layer. The second and third groups are similar to the first, but each contains only 64 filters. Then, we added a max-pooling layer with a pool size of 3 and a stride of 2. After these three groups, we included a flattened layer, two ReLU dense layers (each with 64 nodes), and finally, a Softmax layer.
Our third and last model is a Support Vector Machine (SVM) NIDS. We used the vanilla version with a linear kernel. For all models, we used 80 percent of the selected records from the dataset for training and validation, and the remaining 20 percent for testing.
VI-D The Adversarial Attacks in Our Experiments
The Adversarial Robustness Toolbox (ART) library [38] was used to generate FGSM [15], Projected Gradient Descent (PGD) [39], and Basic Iterative Method (BIM) [40] adversarial attacks for our experiments. We tested the attacks by adding adversarial perturbations to the entire test set.
VI-E Continuous Training
The CSE-CIC-IDS2018 dataset is partitioned by days, and we assumed that the target ML-NIDS model retrained daily. However, in continuous learning environments, there is a risk of catastrophic forgetting. Catastrophic forgetting in machine learning refers to the phenomenon where a neural network or learning model forgets previously learned information when trained on new data or tasks [41].
To address this issue, we merge all data frames from previous days for each new day, shuffle the union set, and select a random sample with a uniform distribution. This approach overcomes catastrophic forgetting by ensuring the model retains its knowledge and adapts to new data without forgetting previously trained data. We repeated this process for three consecutive days to test the impact of continuous training on adversarial attacks. This simple yet effective technique enabled us to retrain the model without performance drops or the need for continuous fine-tuning.
VII Dynamic ML Effect Findings
Section V addressed numerous issues that pose significant obstacles to adversarial attacks against Network Intrusion Detection Systems. However, even if attackers were to overcome these challenges, the implementation of dynamic learning would reduce the effectiveness of the attacks, as seen in this section.
VII-A Attack Scenario
In this section, we will present the results of our experiments that test the effect of the continuous training aspect of the dynamic learning models on adversarial attacks. Our testing scenario, as depicted in Figure 5, assumes that attackers have somehow acquired the knowledge needed to calculate gradients and craft their adversarial attacks, despite the challenges discussed in Section V regarding obtaining information about the target.
Due to the complexity of formulating adversarial attacks against NIDS, there is a time delay (at least one or two days in our experiments) before the attackers initiate the evasion. Meanwhile, the target dynamic ML-NIDS undergoes retraining, rendering the knowledge obtained by the attackers outdated. In our experiments, we aim to test the efficiency of adversarial attacks against ML-based NIDS using up-to-date knowledge of the model (Day n) for evasion versus outdated knowledge after one day of retraining the NIDS (Day n+1) and two days (Day n+2). In other words, we attack the continuously daily trained model using the gradients captured on Day n to attack the model on Day n (up-to-date gradients), Day n+1 (one-day-old gradients), and Day n+2 (two-day-old gradients).
TABLE I: Effect of Continuous Training on FGSM, PGD and BIM Attacks against the ANN, SVM and CNN NIDSs, where Day n represents attacking before re-training, Day n+1 represents attacking one day after re-training, and Day n+2 represents attacking two days after re-training.
(a) Attack: FGSM - Target Model: ANN (b) Attack: FGSM - Target Model: SVM
Accuracy Precision Recall F1 Accuracy Precision Recall F1
Pre-attack 0.997 0.995 1 0.997 Pre-attack 0.998 0.996 1 0.998
Day n 0.756 0.999 0.513 0.677 Day n 0.150 0 0 0
Day n+1 0.938 0.996 0.880 0.934 Day n+1 0.998 0.996 1 0.998
Day n+2 0.966 0.996 0.896 0.966 Day n+2 0.998 0.996 1 0.998
(c) Attack: FGSM - Target Model: CNN (d) Attack: PGD - Target Model: ANN
Accuracy Precision Recall F1 Accuracy Precision Recall F1
Pre-attack 0.997 0.994 1 0.997 Pre-attack 0.997 0.995 1 0.997
Day n 0.842 0.762 0.995 0.863 Day n 0.750 0.999 0.501 0.666
Day n+1 0.957 0.921 0.999 0.959 Day n+1 0.946 0.996 0.896 0.943
Day n+2 0.989 0.981 0.997 0.989 Day n+2 0.929 0.995 0.861 0.924
(e) Attack: PGD - Target Model: SVM (f) Attack: PGD - Target Model: CNN
Accuracy Precision Recall F1 Accuracy Precision Recall F1
Pre-attack 0.998 0.996 1 0.998 Pre-attack 0.997 0.994 1 0.997
Day n 0.150 0 0 0 Day n 0.584 0.551 0.920 0.689
Day n+1 0.998 0.996 1 0.998 Day n+1 0.930 0.900 0.969 0.933
Day n+2 0.998 0.996 1 0.998 Day n+2 0.971 0.961 0.982 0.971
(g) Attack: BIM - Target Model: ANN (h) Attack: BIM - Target Model: SVM
Accuracy Precision Recall F1 Accuracy Precision Recall F1
Pre-attack 0.997 0.995 1 0.997 Pre-attack 0.998 0.996 1 0.998
Day n 0.750 0.999 0.501 0.667 Day n 0.150 0 0 0
Day n+1 0.946 0.996 0.895 0.943 Day n+1 0.998 0.996 1 0.998
Day n+2 0.930 0.995 0.863 0.925 Day n+2 0.998 0.996 1 0.998
(i) Attack: BIM - Target Model: CNN
Accuracy Precision Recall F1
Pre-attack 0.997 0.994 1 0.997
Day n 0.688 0.626 0.939 0.751
Day n+1 0.932 0.902 0.969 0.934
Day n+2 0.979 0.972 0.987 0.979
VII-B Results
In this subsection, we present the results of continuous training on different ML-based NIDSs (ANN, SVM, CNN) and their susceptibility to three different attack techniques. Carlini et al. [42] recommended using strong iterative attacks in addition to FGSM to evaluate adversarial attacks. Therefore, we used FGSM, PGD, and BIM attacks in our experiments. The performance metrics considered were Accuracy, Precision, Recall, and F1-score.
Figures 10 and 11 and Table I illustrate the impact of FGSM, PGD, and BIM attacks on three different NIDSs: Artificial Neural Network, Support Vector Machine, and Convolutional Neural Network. These attacks were conducted on various days. Specifically, ’Day n’ indicates an attack before re-training (with current knowledge of the model), ’Day n+1’ corresponds to an attack one day after re-training (with knowledge that’s one day old), and ’Day n+2’ signifies an attack after two re-training cycles (with knowledge that’s two days old).
Pre-attack, as seen in Table I (a, b, and c), the NIDSs achieved high accuracy, precision, recall, and F1-score, indicating their effectiveness in detecting intrusions. However, after the FGSM attack, the performance dropped significantly for all three models. Among them, the SVM model experienced the highest drop, with its accuracy and F1 decreasing from 0.997 and 0.997 to 0.150 and 0, respectively. On the other hand, the ANN and CNN models had a decrease in accuracy and F1 of 15 to 30 percent. The reduction in accuracy and F1-measure was particularly prominent, indicating that the attacks successfully evaded the NIDSs’ detection. Nevertheless, continuous training improved the NIDSs’ resistance to FGSM attacks. On Day n+1 and Day n+2, we observed a significant recovery in accuracy, precision, recall, and F1-score. As seen in Figures 11a and 10a, after just one day of re-training, all the models recovered an accuracy and F1-score of more than 0.9. Moreover, after one more day of re-training (Day n+2), accuracy and F1-score increased even further, except for the SVM model, where the accuracy and F1 stayed the same at 0.998, which is very high.
Similarly, as shown in Table I (d to i), the results indicate that, like the FGSM attack, the performance of all three models significantly decreased after the PGD and BIM attacks. Among them, the SVM models experienced the most pronounced decline, while the ANN and CNN models exhibited a milder decrease in accuracy and F1. Continuous training also proved to enhance the NIDSs’ resistance to PGD and BIM attacks. On Day n+1 and Day n+2, we observed a substantial recovery in accuracy, precision, recall, and F1-score compared to Day n. As depicted in Figures 11b, 10b, 11c, and 10c, after just one day of re-training, all the models achieved an accuracy and F1-score of over 0.9. For the PGD and BIM attacks, after one additional day of re-training (Day n+2), the accuracy and F1-score of the CNN model further improved. In the case of the SVM model, the accuracy and F1 remained constant at 0.998, a very high level. However, unlike FGSM, the metrics of the ANN model experienced a slight decrease (2 percent) after one more day of re-training (Day n+2) in the cases of BIM and PGD attacks. This minor decrease can be attributed to a slight instance of catastrophic forgetting.
These results suggest that attacks crafted with slightly outdated knowledge of the model might have a lower effect on ML models. That indicates that continuously training the model might introduce an additional hurdle for attackers, as they would constantly need to obtain the updated gradients of the model, which can be a complex task, especially in the NIDS domain. Moreover, it is worth mentioning that this recovery of the model’s measures occurred after just one and two re-training sessions. Furthermore, these results were achieved without adversarial training, meaning that the re-training data didn’t contain adversarial samples [20].
Figure 12: Decision Boundary Evolution using t-SNE
Figure 13: Data Distribution Evolution using t-SNE, wher color intensity encodes data density, with lighter areas representing higher concentrations of points.
On another note, although the results between the three attacks used in our experiments, FGSM, PGD, and BIM, are close, they are different attacks with distinct generation algorithms and dissimilar generation times, as seen in Table II. FGSM is a one-shot attack, whereas PGD and BIM are iterative attacks. This suggests that the impact of continuous training on adversarial attacks may depend more on the ML models than on the attack methods themselves, which is why we have included all these results in the paper. Also, these results demonstrate that the impact of continuous training remains consistent across different attacks. Furthermore, the results reveal that various ML models exhibit different responses to adversarial attacks and varying recovery rates after retraining. These similarities and dissimilarities warrant further investigation in the future.
TABLE II: Generation of Adversarial Attacks Average Time
Attack Model ANN SVM CNN
FGSM 7.2 sec 140.5 sec 42.3 sec
PGD 51.2 sec 2137.0 sec 338.9 sec
BIM 40.2 sec >30 min 305.3 sec
VII-C Interpretation of the Results
Because numerical results alone don’t tell the entire story, and to gain an intuitive understanding of the data and results, visualizations are necessary. To achieve this, we combine the visualization of Voronoi-based representations of multidimensional decision boundaries in scatterplots with the visualization of distances to the multidimensional boundary of data elements [43]. Data elements are reduced from 50 dimensions to 2 dimensions using the t-SNE (t-distributed Stochastic Neighbor Embedding) algorithm [44]. In other words, we show the evolution of the optimal multidimensional decision boundaries in 2D after every re-training session, as seen in Figure 12. Moreover, we also visualize the evolution of the data distribution after every re-training, as shown in Figure 13. As expected, the decision boundaries and training data distribution are always changing and evolving, as seen in Figures 12 and 13.
On the other hand, to generate an adversarial example, small perturbations are added to alter the class of a sample. In other words, it attempts to move the sample beyond a nearby decision boundary [45]. As illustrated in Figure 14, perturbations (arrow) are added to a sample (blue circle) to push it into the other class, creating an adversarial example (yellow circle). It has been observed that adversarial samples are typically located near decision boundaries [46] due to the similarity constraint in adversarial attacks [24]. This constraint aims to ensure that the generated adversarial example remains similar to the original input to a certain degree while still fooling the target model. In simple terms, it limits how much the adversarial example can deviate from the original input while still being effective in fooling the model.
So re-training changes the model, including the decision boundary and gradients, and adversarial examples tend to be near this constantly evolving boundary rather than deep within the wrong class. Therefore, attacks formulated with outdated knowledge of the model might have a reduced effect on machine learning models even without adversarial training.
VIII Discussion and Lessons Learned
This section discusses several lessons learned and insights gathered from our research.
VIII-A Dynamic Learning
  • Dynamic ML research is scarce in the machine learning-based NIDS domain. This scarcity needs attention because it is expected that concept drifts will happen in the NIDS space similar to other research domains [7]
  • The dynamic learning effect on adversarial attacks depends on numerous factors such as the model used, re-learning rate, re-learning data, etc. Our work just shows that it could reduce the effect of adversarial attacks on attacked models.
  • Systems that adapt to changes in legitimate behavior over time, such as evolving machine learning approaches, may be vulnerable to manipulation by intruders who gradually introduce their own malicious behavior into the baseline. This is known as session creep [47].
  • In continuous learning environments, a critical challenge is the potential occurrence of catastrophic forgetting, a phenomenon where machine learning models forget previously learned information when trained on new data or tasks [41]. To ensure the successful deployment of such models in real-world conditions, effective strategies for handling catastrophic forgetting must be carefully implemented.
VIII-B Network Security Datasets
  • Obtaining Labeled datasets to train even static ML-NIDS is very tough and expensive [48]. Unfortunately, dynamic ML models need more labeled datasets for continuous training, and acquiring these labeled datasets is even harder. This topic needs more research from network security and ML communities.
  • Although CSE-CIC-IDS-2018 is widely regarded as a reliable/up-to-date NIDS dataset [11], it has been discovered that it has several issues with feature generation and labeling [34]. This re-emphasises the issue that the ML and NIDS research communities should prioritize the development of more robust datasets.
VIII-C Adversarial attacks in network security vs computer vision
  • Crafting adversarial attacks in network/cybersecurity is more complex than in domains like computer vision [31]. Unlike subtly altering pixel values in image classification, adding perturbations to network packets or PE files can disrupt their network or malicious functionality. White-box attacks require extensive knowledge of the targeted ML model. This is often impractical in many domains like computer vision but even more challenging in network security due to heightened security measures [31]. Black-box attacks assume querying the model for gradients, but this may be unfeasible in Network Intrusion Detection Systems (NIDS) contexts.
  • On the other hand, if attackers manage to overcome all obstacles and create a practical problem space adversarial attack in the network/cybersecurity domain, the perturbed network elements and features will become more diverse and heterogeneous. This complexity makes detecting and defending against problem space adversarial attacks in the network/cybersecurity domain more challenging than in other domains, such as computer vision [19].
Figure 14: Adversarial Attacks Visualization
IX Conclusion and Future Work
In this paper, we extended the research discussion on the practicality of adversarial attacks against NIDS by identifying the questionable practicality prerequisites required to carry out evasion adversarial attacks against NIDS using an attack tree and by presenting a taxonomy of practicality issues related to ML-based NIDS adversarial attacks. Our attack tree and taxonomy highlight several factors that could make numerous researched adversarial attacks impractical against real-world ML-based NIDSs. It is important to note that our work does not claim that adversarial attacks won’t harm ML-based NIDSs; rather, we find that the gap between research and real-world practicality is wide and deserves to be addressed. Apruzzese et al. [3] argue that real-world attackers typically do not rely on gradients or conventional adversarial attacks techniques; instead, they employ alternative simple yet effective approaches to evade ML models. If this is the case, then the current direction of ML security research needs a significant shift to address real-world challenges adequately. We also explored the impact of adversarial attacks on dynamic ML-based NIDS, finding that continuous re-training, even without adversarial training, may limit the effect of such attacks. In future research, We plan to expand our attack taxonomy and apply it to network security applications beyond NIDS (e.g., spam detection). We will also explore how more sophisticated dynamic and continuous learning techniques will perform against adversarial attacks.
X Acknowledgement
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through the NSERC Discovery Grant program.